OpenStudy (anonymous):

find all local max,min and/or saddle point of f(x,y)=4+x^3+y^-3xy

OpenStudy (anonymous):

would you plz write the function using equation

OpenStudy (amistre64):

hmm, something to do with:\[f_{xy}f_{yx}-(f_{xx})^2\] or did i recall that incorrectly?

OpenStudy (amistre64):

almost had it :) \[f_{xx}f_{yy}-(f_{xy})^2\]

OpenStudy (anonymous):

\[4+x ^{3}+y ^{3}-3xy\]

OpenStudy (anonymous):

it fxx.fyy-(fxy)^2

OpenStudy (anonymous):

i jst want to verify the answer for it !

OpenStudy (turingtest):

what answer did you get?

OpenStudy (anonymous):

\[fx = 2x-3y \] \[fy= 2y-3x\] on solving both equations the point is (0,0) \[fxx = 2\] , \[fyy= 2\] \[fxy = \] \[D = -5 <0 saddle point\]

OpenStudy (turingtest):

that's not quite right I don't think

OpenStudy (anonymous):

fxy = -3

OpenStudy (turingtest):

\[f(x,y)=4+x^3+y^3-3xy\]\[f_x=3x^2-3y=0\implies y=x^2\]\[f_y=3y^2-3x=3x^4-3x=3x^3(x-1)=0\]\[ x=\{0,1\},y=\{0,1\}\]

OpenStudy (anonymous):

its x^3 so i have (0,0) ,(1,1)and a (-1,1) i am not sure if (-1,1) should be there oor not

OpenStudy (anonymous):

@Living_dreams I solved what you mentioned in your comment

OpenStudy (turingtest):

there is no x=-1 solution, I don't see how you get that

OpenStudy (turingtest):

@Avva I don't see how you get your fx and fy

OpenStudy (anonymous):

Ooh I saw the powers on X ^ y 2 not 3

OpenStudy (anonymous):

@Avva thanks for ur help but i was wonder abt the max and min points! n\ @TuringTest : 3y^2-3x=0 i sub y=x^2 from other equation so it will be 3x(x^3-1)=0 therefore you have x =0,1,-1 i am not sure abt -1 though !! tht was my question

OpenStudy (turingtest):

x^3=1 does not have x=-1 as a solution since (-1)^3=-1, not 1

OpenStudy (turingtest):

I did screw up my factoring earlier though, you were right about that :P

OpenStudy (turingtest):

3x(x^3-1)=0 3x=0 -> x=0 x^3=1 -> x=1

OpenStudy (anonymous):

@TuringTest that alrite!! :) so now my question is x^3=1 so will i hv x=-1 if i solve for x ?

OpenStudy (turingtest):

no, again, because \((-1)^3\neq1\) so \(x^3=1\) has only one real solution, x=1 x=-1 is NOT a solution

OpenStudy (anonymous):

the other two roots are complex not real

OpenStudy (anonymous):

@TuringTest oh ok thanks a lot for ur help!! so the final ans will be (0,0 ) is a saddle point and (1,1) will be a local minimum !

OpenStudy (turingtest):

exactly :)

OpenStudy (anonymous):

@Avva :thanks for ur help !! :)

OpenStudy (anonymous):

@Living_dreams URW :) sorry I didn't get that the powers are 3 at first

OpenStudy (anonymous):

it okay !!i knw it get hard to tell the difference between 2 and 3 if we type it in equation form !!